OpenStudy (anonymous):

Solve Equation please last one √5x-1 = √4x+9

OpenStudy (saifoo.khan):

\[\sqrt{ 5x-1 }= \sqrt{4x+9}\]Take square on both sides..

Parth (parthkohli):

@saifoo.khan You go.. :D

OpenStudy (saifoo.khan):

5x - 1 = 4x + 9 5x - 4x = 9 + 1 x = 10

OpenStudy (saifoo.khan):

@ParthKohli . :)

Parth (parthkohli):

Don't give me achieved medals that I didn't achieve... take that backk!!!

OpenStudy (anonymous):

Make me ;3